Chair: Piet Martens (SOC) Notes: Laura Allen (SOC) Present:Ed Schmahl (UMd), Jacques Charra (GOLF), Rosine Lallement (SWAN), Leonard Strachan (UVCS), Andrzej Fludra (CDS), Bernhard Fleck (ESA), Udo Schuehle (SUMER), Chris StCyr (LASCO), Barbara Thompson (EIT), Julia Saba (MDI), Rock Bush (MDI), Craig DeForest (MDI), Luis Sanchez (ESA), Vicente Domingo (Project Scientist)

1. Review of action items:

97_2.1: Working group to be formed for the SOHO archive, in particular the issue of linking the campaign database to the experiment catalogs: A congnoscente from each instrument team, SOC, Luis, Macwan, Liyun, Dominic. CDS -- Bill Thompson SUMER -- Ingolf Dammasch UVCS -- Angela Ciaravella EIT -- Joe Gurman, Jeff Newmark LASCO -- TBC, Russ Howard MDI -- Phil Scherrer or Rock Bush Luis Sanchez will be in charge and call meetings. STATUS: On-going. Meeting to be held on Monday, March 24. 97_2.2 Action on EIT to provide SOC information concerning desired EIT offpointing in the next 6 months. STATUS: Open. SOCs have received no information concerning this.

2. Reengineering

A handout was distributed. Someone of the re-engineering team will report at the SWT in April.

3. Boundary conditions for the coming month

- MDI Continuous: Starts April 13, until July 13 - SVM Memory Dump, 6 hours, April 1 - Momentum Management and Station Keeping, April 11

4. Preview of future months (including Medoc campaign)

- See Monthly Calendar - MEDOC campaign, proposed for May 26 - June 8, possibly one week before or after as well. CDS, and SUMER will to participate. No response yet from EIT and LASCO. MDI will of course be in continuous contact during this time and require little in the way of planning, and will not be participating in joint observations. Basic idea is that all the planning will be done from MEDOC. CDS planning will be at MEDOC but the commanding will be done from the EOF. SUMER planning and commanding will be from MEDOC. The observing campaigns from Medoc will include collrborations with Themis, prominence observations with Meudon and Pic-du-Midi, and JOP056, "The Interstellar Focusing Cone", led by SWAN. The final decision will be taken by the SWT in its April meeting.

5. Priorities for the coming month

a) Individual instrument plans MDI -- MDI showed a plot of pointing adjustments. Next week will be performing full flat field measurements, and then moving MDI legs to re-center images. MDI sees a drift less than an MDI pixel, and can see the effect of sunspots as they pass by the photodiodes. MDI presented an update of their upcoming 3-day continuous contacts: -- Full disk 60 sec cropped velocity and split magetogram -- Two full disk 30 sec velocity with sub-limb cropping (480 pixel crop radius) -- Full disk 30 sec cropped velocity and HR 1024 x 640 velocity For the 2 month continous: full disk doppler and magnetograms last two weeks, high cadence program of 30 second dopplergrams Details of the MDI high resolution field-of-view can be found at: MDI will have no planners during the contiuous contact, since all it will be providing for collaborations will be the full-disk magnetograms. MDI needs to know deadline for OCD for continuous operations. LASCO -- LASCO planning a new software upload in late April. LASCO will then have a new image compression algorithm using wavelets. The software update will also fix EIT's sector motor problems. After the 2 weeks of high rate telemetry, there will be a push to do more spectroscopy and observe the K corona better. LASCO also aims to fix Fabry Perot wavelength position. SUMER - Until 30 March will be observing. Then giving 2 weeks of science telemetry to LASCO/EIT and closing the door. SUMER will make sure one person is on call in area when their door is closed. SUMER will be resuming observations ~ 14 April at the latest. They are currently installing workstations at MEDOC. Around April 28, SUMER will be testing and sending science commands for 4 hours per day for 2-3 days. They will be testing commanding from both workstations and PCs at MEDOC. Later in the year, SUMER may switch to some operations from MEDOC. CDS -- Andrzej Fludra thanks EIT, MDI, SUMER, for JOP021 support over the last month. CDS plans active region studies, study of plumes above limb, macrospicule searches, and opacity studies in the coming month. - 3 April - 5 May, JOP059, St Andrews Loop Campaign. - 19 April -- Woods/Hassler rocket flight (now delayed unitl May 9) - 21 April - 27 April, (Week 17) Helium Enhancement Study, which will be submitted as a JOP Also during that week, will run JOP054, Scaling Laws of Coronal Loops UVCS -- The lead observers and science programs for April 1997 are the following: 05 May - 11 May John Kohl Global dynamics of the extended corona and CME Watch with LASCO 28 Apr - 04 May Ester Antonucci Global dynamics of the extended corona and CME Watch with LASCO 21 Apr - 27 Apr Shadia Habbal Coronal filling factor and outflows in streamers 14 Apr - 20 Apr Steven Cranmer Latitudinal Dependance of Nonradial Wind Expansion; Coronal Holes at large heliographic heights. 07 Apr - 13 Apr Larry Gardner UVCS Calibration; Intercal 01 31 Mar - 06 Apr Carlo Benna Polar Plume Studies and CME Watch with LASCO SWAN -- SWAN will observe comets, and proposes a run of JOP056 for the first week of June -- Helium detection with UVCS and SUMER. Eric Quemerais will support/coordinate this JOP. GOLF -- Nominal. Observations continue. EIT -- Finished a month of 304 angstrom CME watch every 17 minutes. During the last week's momentum management obtained 304 angstrom images every 6 minutes. The next month will perform CME watch observations at 195 angstroms, with LASCO C2 coronagraph. EIT will stay with the 195 bandpass, especially during the high rate when they are in submode #6. b) Intercalibration activities Udo Schuehle would like run Intercal 001 during the second half of April. UVCS wanted second week of April; however, Udo claimed this was too early. Udo picked Thursday, 24 April at 16:00 UT as a date to run Intercal001. c) Joint observing plans A full schedule for April was layed out. See the SOHO Calendar for the latest version. For a link to the Siberian Solar Radio Telescope, which will observe togther with SOHO and Nobeyama at the end of April, see:

6. JOPs status

- JOP-leaders' reports Starting next week, JOP leaders will see incessant e-mail to update their JOP information in the campaign catalogue. - Campaign and Events Catalogs Campaign Catalogue now complete until August 31 of last year. After that all campaigns are in catalog, but information per campaign is not always complete.

7. New JOPs

- JOP060, VLA-SOHO Observations of Transient Sources in the Solar Corona, Kenneth Lang and Rob Willson. Also may involve Big Bear Observatory. Basically, a re-run of last year's observations with the VLA. - JOP061, Brightening Flows and Waves, Isabelle Ruedi.

8 . AOB

- Conferences and Workshops: * Fifth SOHO Workshop, Oslo, June 17-20. Second announcement will be send out next week. * Second SOHO - Yohkoh Worshop: will be at MEDOC around the fall equinox. Bob Bentley, Jean-Claude Vial and Helen Mason organizers. Topic will be Active Regions * ESLAB Symposium -- late September at ESTEC. * ISTP January 6 Event Worskshop --- will be held April 8-9 to discuss January 6th event and similar occurrences (such as February 7th). People are encouraged to post their results on the Web first so that attendees have a chance to look at the results before attending the workshop. - Bernhard Fleck to receive a bottle of Californian wine from Craig DeForest for winning a bet made concerning the availability of the calibrated data from some of the spectral instruments at this time. The optimist won -- the data are available! - Luis Sanchez reported that the DEFINITIVE orbit files prior to November 1996 (in which CDHF did not rotate the coordinates of the velocity vector properly), have finished being reprocesssed. We now have a consistent set of data. - Luis Sanchez also reported that due to storage space considerations, we are forced to move the entire SOHO public tree to a different Network Appliance. Instead of soho-arch00, this area will live now on soho-arch03 (that will have enough capacity even for an extended mission!). ADJOURNED, 2:45 pm